OpenStudy (anonymous):

What value of a makes u and v orthogonal, if u = and v = ? -8 2 4 8 16

OpenStudy (unklerhaukus):

Th dot product of orthogonal vectors, should be zero

OpenStudy (usukidoll):

orthogonal means 0... usually we use the dot product to figure out whether or not it's orthogonal

OpenStudy (usukidoll):

you have to plug in those values and use dot product and see if one of your choices produces 0

OpenStudy (usukidoll):

try 4 = a

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I get a 16

OpenStudy (anonymous):

or a -8

OpenStudy (usukidoll):

what about that -2 x 8 ?

OpenStudy (anonymous):

a -16

OpenStudy (anonymous):

(a-4) (a+4)

OpenStudy (usukidoll):

ok wait wait wait... dot product = u1v1 +u2v2 where u = <u1,u2> and v= <v1,v2>

OpenStudy (usukidoll):

dot product < - we're multiplying not adding vectors

OpenStudy (usukidoll):

let's say if u = <a, -2 > and v= <a, 8> and a =4.. then u = <4,-2> and v < 4.8> now taking dot product I get (4)(4)+(-2)(8) = 16-16=0

OpenStudy (anonymous):

so a*a and 8*-2

OpenStudy (michele_laino):

another way, for solving your exercise, is: using the formula for scalar product, we get: a^2-16=0

OpenStudy (usukidoll):

yeah .. do you know the dot product formula?

OpenStudy (usukidoll):

true but that is not how I learned it.

OpenStudy (usukidoll):

btw a = 4, -4

OpenStudy (usukidoll):

I rather have it u dot v = u1v1+u2v2....unvn

OpenStudy (usukidoll):

even he has it like that http://tutorial.math.lamar.edu/Classes/CalcII/DotProduct.aspx if let a = 4 and use dot product you will have 0

OpenStudy (anonymous):

<4,-2>, <4,8>

OpenStudy (usukidoll):

yes now use the dot product formula ^^

OpenStudy (usukidoll):

u dot v = u1v1+u2v2+...unvn where u = < u1, u2....un> and v =<v1,v2... vn>

OpenStudy (anonymous):

=2

OpenStudy (usukidoll):

whattttttttttt

OpenStudy (anonymous):

maybe I am looking at the formula wrong

OpenStudy (usukidoll):

where is your u1 and v1?

OpenStudy (anonymous):

4,4

OpenStudy (usukidoll):

yes so what is 4 x 4 ?

OpenStudy (anonymous):

16

OpenStudy (usukidoll):

yes so what is u2 and v2?

OpenStudy (anonymous):

-2,8

OpenStudy (anonymous):

the answer is 4

OpenStudy (anonymous):

I was reading wrong my fault

OpenStudy (usukidoll):

yes a =4 because 4 x 4 + (-2) x 8 = 16-16=0

OpenStudy (anonymous):

yea orthogonal is 0

OpenStudy (anonymous):

so yea 4

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

That is probably why I was adding the vectors

OpenStudy (anonymous):

my fault :(

OpenStudy (usukidoll):

it's ok.. as long as you know that dividing vectors is impossible.